OpenStudy (anonymous):

sqrt x +1= sqrt x +7= solve and check answer

geerky42 (geerky42):

try to square both sides.

OpenStudy (xapproachesinfinity):

Yes square both side to get x+1=x+7 and solve

OpenStudy (xapproachesinfinity):

that if square cover x+1 and x+7

OpenStudy (xapproachesinfinity):

that is*

OpenStudy (aum):

You have to make it clear where the square root stops with x + 1. Are both x and +1 within the square root or just x? Same thing with x + 7. Are both x and +7 within the square root or just x?

OpenStudy (aum):

Use parenthesis to make it clear.

OpenStudy (anonymous):

x+1 within sqrt = x+7 within sqrt

OpenStudy (aum):

\[ \sqrt{x+1} = \sqrt{x+7}~~ ? \]

OpenStudy (anonymous):

yes exactly

OpenStudy (aum):

There is NO solution to that.

OpenStudy (aum):

There is no value of x to which you can add a 1 OR a 7, take the square root and end up with the same value.

OpenStudy (anonymous):

there is a very short sqrt over x+1 then a long sqrt over x+7

OpenStudy (aum):

\[ \sqrt{x} + 1 = \sqrt{x+7} ~~~? \\ \]Don't say exactly unless you are sure.

OpenStudy (anonymous):

yes I am looking at the paper.

OpenStudy (aum):

You can post a screenshot of the question if you can.

OpenStudy (anonymous):

don't know how but you have it as written like it is on the paper. I only have a few minutes to answer the next few question.

OpenStudy (aum):

\[ \sqrt{x} + 1 = \sqrt{x+7} \\ \text{Square both sides:} \\ x + 1 + 2\sqrt{x} = x + 7 \\ 2\sqrt{x} = 6 \\ \text{Square both sides:} \\ 4x = 36 \\ x = 9 \]

OpenStudy (aum):

substitute back and make sure x = 9 works.

OpenStudy (anonymous):

thank you Aum. I have one more for you. Long sq rt Y-9=short sqrt Y-1

OpenStudy (aum):

Use the same method shown above.